OpenStudy (anonymous):

Can any one teach me all the graphs in motion in a straight line chapter , including the way the questions will be asked from the graphs and extras , please ?

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

what do u want to know....

OpenStudy (anonymous):

everything .. the position-time graph at rest , uniform motion , non uniform motion dist-time graph .. almost every graphs !

OpenStudy (anonymous):

OK now I am free......

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

do u know what a uniform motion means

OpenStudy (anonymous):

yeah motion in a straight line ?

OpenStudy (anonymous):

covering equal distance in equal intervals of time .

OpenStudy (anonymous):

I dont think so.....

OpenStudy (anonymous):

then ?

OpenStudy (anonymous):

yeah the second one is correct

OpenStudy (anonymous):

so first one is not right ?

OpenStudy (anonymous):

yeah......

OpenStudy (anonymous):

okay !

OpenStudy (anonymous):

I think u only need linear uniform motion......I mean uniform motion in a straight line

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yeah !

OpenStudy (anonymous):

so.... in a linear uniform motion.... accleration is 0 so we have only one formula.... and that is s=ut

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

okay !

OpenStudy (anonymous):

Now, lets talk about non uniform linear motion.....

OpenStudy (anonymous):

okay ..

OpenStudy (anonymous):

here the velocity of the body changes so we will have accleration.. right?

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

in a non uniform linear motion . velocity changes .... right?

OpenStudy (anonymous):

yeah !

OpenStudy (anonymous):

so...... if velocity changes there is always accleration right?

OpenStudy (anonymous):

yeah okay , got it :)

OpenStudy (anonymous):

So we have some equation of motions..... to descibe non uniform linear motion

OpenStudy (anonymous):

I think u know them........

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

v = u+at s = ut + 1/2at^2 v^2 - u^2 = 2as these ones ?

OpenStudy (anonymous):

yep.......

OpenStudy (anonymous):

okay !

OpenStudy (anonymous):

now, while using these formula u must be very careful

OpenStudy (anonymous):

because velocity, and accleration are vector quantites

OpenStudy (anonymous):

so their sign differs with respect to the direction

OpenStudy (anonymous):

so what's the problem ?

OpenStudy (anonymous):

oh !

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

okay .. like ?

OpenStudy (anonymous):

if u take